# PLD: A Choice-Theoretic List-Wise Knowledge Distillation

### v2

@@ 4.2 Confidence-Weighted Likelihood Yielding PLD
In the standard PL model, each selection step is weighted equally-i.e., every position contributes identically to the likelihood. Position-aware ListMLE (P-ListMLE) lan2014position addresses this by introducing a fixed, strictly decreasing weight sequence-e.g. $\alpha_{k}=2^{\,C-k}-1,$
so that errors in top-ranked positions are penalized more heavily than those further down. However, in multiclass classification-where only the top-1 decision ultimately matters-fixed, hand-crafted weight schedules are both hyperparameter-sensitive and fail to reflect the inherently greater importance of the first selection compared to subsequent ones. Observing that a pretrained teacher naturally assigns far greater confidence to the correct class than to the others, we therefore parameterize the weight at step $k$ by the teacher’s softmax probability: $\alpha_{k}=q^{T}_{\pi^{*}_{k}}$
This choice ensures the loss automatically emphasizes the top-1 selection when the teacher is confident yet relaxes its focus when the teacher’s output distribution is more uniform. Furthermore, our data-driven weights recover several well-known ranking surrogates as special cases: setting $\alpha=(1,0,\dots,0)$ recovers standard cross-entropy; choosing a uniform distribution $\alpha=(\tfrac{1}{C},\dots,\tfrac{1}{C})$ yields a ListMLE-like objective; and any fixed, strictly decreasing sequence reproduces the P-ListMLE surrogate.
@@ 5 Experiments
All experiments are conducted on ImageNet-1K deng2009imagenet following the recipe wightman2021resnet with minimal data augmentation. We train for 100 epochs with an effective batch size of 2048 images (256 per GPU across eight NVIDIA A100 SXM4 80 GB accelerators) using the LAMByou2019large optimizer, an initial learning rate of $5\times 10^{-3}$ decayed by a cosine schedule and linearly warmed up over the first 5 epochs, and weight decay fixed at 0.02. Under this setup, 100 epochs require approximately 6 hours for convolutional backbones and 8 hours for Vision Transformers. To focus on the impact of our loss formulations, we apply only random resized crops, horizontal flips, and per-channel normalization to ImageNet statistics.
@@
For validation we adopt the “A-recipe” from wightman2021resnet . Specifically, we set the test resolution $r=224$ and test crop ratio $\rho=0.95$, then $\text{resize}_{\min}=\bigl{\lceil}r/\rho\bigr{\rceil}\approx 236,$ apply a bicubic resize of the shorter side to $\text{resize}_{\min}$, followed by a center crop of size $r\times r$. Finally, we normalize by the same mean and standard deviation used during training.
5.1 Baseline Configuration and Ablation
We evaluate four losses-CE, KD hinton2015distilling , DIST huang2022knowledge , and PLD-using a ResNet-50 student and ResNet-152 teacher He2015 under identical training. Each objective is a drop-in replacement of the logit-based term and incurs comparable cost, so KD (divergence-based) and DIST (correlation-based) serve as natural baselines against our ranking-based PLD. We sweep each method’s key hyperparameters ($\alpha$ for KD, $\beta,\gamma$ for DIST, $\tau_{T}$ for PLD) to find optimal settings. Figure 1(a) shows the effect of varying the CE mixing weight: gradually reducing $\alpha$ improves both KD and DIST up to an optimal region near $\alpha\approx 0.1$, but fully dropping CE ($\alpha=0$) degrades performance. Table 1 reports results with and without the CE term (full ablation in Appendix B), and Table 2 gives PLD-specific sweeps.
Table 1: Top-1 and Top-5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “–” indicates not applicable. CE: standard cross-entropy. LS:szegedy2016rethinking cross-entropy with label smoothing ($\epsilon=0.1$). KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$. DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.Top-5 Acc.%
Teacher (ResNet-152) 79.61
CE 71.35
LS ($\epsilon=0.1$) 73.92
KD 0.00 2.00 75.92 92.82
0.10 2.00 76.80 93.16
DIST 0.00 0.50 0.50 1.00 76.47 93.19
0.10 0.45 0.45 1.00 76.60 93.30
Table 2: Top-1 and Top-5 accuracy for PLD variants (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). “–” indicates not applicable.
Variant $\tau$ Top-1 Acc.Top-5 Acc.%
PLD 0.50 76.06 92.39
1.00 77.30 93.28
1.50 77.17 92.74
2.00 76.46 92.34
4.00 75.22 91.34
PLD–ListMLE 74.11 90.60
PLD–pListMLE 76.60 93.23
5.1.1 Ablation of PLD
As PLD weights each position by $\alpha_{k}\;=\;\frac{\exp\bigl{(}t_{\pi^{*}_{k}}/\tau_{T}\bigr{)}}{\sum_{j=1}^{C}\exp\bigl{(}t_{\pi^{*}_{j}}/\tau_{T}\bigr{)}}$, we sweep $\tau_{T}\in\{0.5,1.0,1.5,2.0,4.0\}$ to assess sensitivity: lower $\tau_{T}$ sharpens the distribution, while higher $\tau_{T}$ flattens it. We also test two special cases: Uniform (ListMLE): $\alpha_{k}=1/C$. Position-aware (p-ListMLE): $\alpha_{k}=q^{T}_{\pi^{*}_{k}}/\sum_{j}q^{T}_{\pi^{*}_{j}}$. As Table 2 shows, the unsoftened PLD ($\tau_{T}=1$) consistently achieves the highest Top-1/Top-5 accuracies, with uniform ListMLE trailing by 2-3 pp. Overall, the best results occur at $\tau_{T}=1.0$ (no softening) or with only slight perturbations ($1.0\pm\epsilon$), demonstrating that PLD is robust to the exact choice of $\alpha$ distribution and does not require aggressive sharpening or over-softening.
5.2 Key Experiments
To evaluate PLD’s robustness and generality, we conduct three complementary studies using hyperparameters identified in our baseline experiments (KD: $\alpha=0.1,\tau=2$; DIST: $\beta=0.45,\gamma=0.45,\tau=1$; PLD: $\tau_{T}=1$). First, in a homogeneous setting, we pair larger teachers with smaller students within the same architecture family to isolate the impact of model capacity on distillation. Second, in a heterogeneous setup, we fix a single student (ResNet-50) and vary teacher architectures across convolutional and transformer families to test PLD under diverse teacher distributions. Finally, we return to the homogeneous scenario with extended pretraining, pretraining a MobileNet-v4qin2024mobilenetv4 Medium student from a MobileNet-v4 Large teacher for 300 epochs to verify that PLD’s gains persist under longer optimization.
5.2.1 Distillation Across Homogeneous Architectures
As illustrated in Figure 2(a), we evaluate three backbone families-ResNetHe2015 , MobileNet-v4qin2024mobilenetv4 , and Vision Transformersdosovitskiy2020image -in a homogeneous setting. Across convolutional models (ResNet, MobileNet-v4), PLD outperforms KD by 0.5-2.7 pp and DIST by 0.2-0.7 pp. On Vision Transformers (ViT), PLD yields +0.72 pp over DIST and +0.30 pp over KD. Averaged across all pairs, PLD improves Top-1 accuracy by +0.42 pp versus DIST and +1.04 pp versus KD. Table 3 reports the full Top-1 results for KD, DIST, and PLD.
5.2.2 Distillation Across Diverse Teachers
One of the key benefits of logit-based distillation is its ability to leverage teacher models of different architectures, provided their output logits share the same dimensionality as the student. Figure 2(b) shows the heterogeneous setting: we fix the student to ResNet-50 ($\approx 25.56$ M parameters) and distill from a range of convolutional and transformer teachers. Across all teachers, PLD consistently outperforms both KD and DIST: its $\Delta_{\mathrm{DIST}}$ gains range from +0.34 pp (MobileNet-v4 Hybrid) to +0.70 pp (ResNet-152), averaging +0.48 pp, while $\Delta_{\mathrm{KD}}$ gains span +0.49 pp to +1.55 pp, with an average of +1.09 pp. As teacher capacity increases, KD’s benefit declines and DIST shows a modest upward trend; PLD not only follows this trend but further widens the margin over DIST for larger teachers. The largest gap between DIST and PLD appears in the homogeneous ResNet-50←ResNet-152 pairing. Table 4 reports each teacher’s standalone Top-1 accuracy, parameter count, and student Top-1 results under KD, DIST, and PLD, along with $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$.
@@ 5.2.2 Distillation Across Diverse Teachers
Table 3: Top-1 accuracy (%) and student model size across student–teacher pairs on ImageNet-1K. “Teacher Acc” is the teacher’s standalone Top-1; $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ are PLD’s gains over DIST and KD.
@@
Table 4: Top-1 % accuracy of ResNet-50 distilled from diverse teachers on ImageNet-1K. “Params” is the teacher’s parameter count in millions. $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ report PLD’s gains over DIST and KD, respectively.
@@
5.2.3 Consistency under Extended Training
As observed in our homogeneous experiments5.2.1 -where KD sometimes outperforms DIST and vice versa-we sought to confirm whether PLD consistently surpasses both. To this end, we independently trained a MobileNet-v4 Medium student from a MobileNet-v4 Large teacher for 100 and 300 epochs using identical hyperparameters. Table 5 reports Top-1 accuracies on ImageNet-1K, and Figure 1(b) provides a detailed comparison. After 100 epochs, PLD achieves 75.72% (vs. DIST +0.32 pp, KD +0.86 pp), rising to 76.94% at 300 epochs (vs. DIST +0.38 pp, KD +0.80 pp). We observe that the distillation gains with PLD (+1.22 pp) are nearly twice those of standard pretraining (+0.67 pp), and that this scaling is comparable across KD (+1.28 pp) and DIST (+1.16 pp). This demonstrates that PLD not only preserves but slightly widens its distillation advantage under extended training.
Table 5: Top-1 accuracy (%) of the MobileNet-v4 Medium student distilled from a MobileNet-v4 Large teacher after 100 and 300 training epochs. The “$\Delta$” row shows the improvement from 100 to 300 epochs.
@@ 6 Conclusion and Limitation
In this work, we introduced Plackett-Luce Distillation (PLD), a unified choice-theoretic framework for logit-based knowledge distillation. Empirically, PLD consistently outperforms both classical KL-based distillation (KD) and recent correlation-based methods (DIST) across a wide range of student-teacher pairs-homogeneous and heterogeneous architectures alike-and remains robust under extended training. These results demonstrate that transferring the teacher’s structured preferences yields stronger and more stable student performance than marginal- or correlation-based objectives.
While PLD applies broadly to any setting in which student and teacher share the same logit dimensionality, it assumes fully aligned class vocabularies and does not directly support mismatched label sets or incremental class addition. Moreover, PLD requires a sorting operation over the $C$ output logits to extract the teacher-optimal permutation, incurring an $O(C\log C)$ complexity per example; although efficient in distributed implementations, this contrasts with the $O(C)$ cost of standard KD or DIST. Finally, as with other distillation methods, PLD’s benefits scale with teacher confidence and may diminish when the teacher’s softmax distribution is nearly uniform.
PLD’s choice-theoretic foundation suggests several promising directions. First, adaptive or curriculum-driven weighting schemes could modulate the ranking loss according to sample difficulty. Second, extending PLD to other domains-such as sequence modeling, reinforcement learning, or multitask settings-may yield similar gains. We hope this work inspires continued exploration of ranking-based objectives for principled and effective model compression.

### v3

@@ 4.2 Confidence-Weighted Likelihood Yielding PLD
In the standard PL model, each selection step is weighted equally-i.e., every position contributes identically to the likelihood. Position-aware ListMLE (P-ListMLE) lan2014position addresses this by introducing a fixed, strictly decreasing weight sequence, such as $\alpha_{k}=2^{\,C-k}-1.$
This scheme penalizes errors at top ranks more than those lower down. However, in multiclass classification-where only the top-1 decision ultimately matters-fixed, hand-crafted weight schedules are both hyperparameter-sensitive and fail to reflect the inherently greater importance of the first selection compared to subsequent ones. Observing that a pretrained teacher naturally assigns far greater confidence to the correct class than to the others, we therefore parameterize the weight at step $k$ by the teacher’s softmax probability: $\alpha_{k}=q^{T}_{\pi^{*}_{k}}$
This choice ensures the loss automatically emphasizes the top-1 selection when the teacher is confident yet relaxes its focus when the teacher’s output distribution is more uniform. Furthermore, our data-driven weights recover several known ranking surrogates as special cases. Setting $\alpha=(1,0,\dots,0)$ gives the standard cross-entropy. A uniform $\alpha=(\tfrac{1}{C},\dots,\tfrac{1}{C})$ produces a ListMLE-like objective, and any fixed decreasing sequence yields the P-ListMLE surrogate.
@@ 5 Experiments
We evaluate PLD on three representative visual recognition datasets: CIFAR-100 cifar100 , ImageNet-1K deng2009imagenet , and MS-COCO coco . These datasets cover small- and large-scale image classification and general object detection.
We first benchmark PLD on CIFAR-100 using traditional convolutional architectures and a unified training recipe. We then assess scalability on ImageNet-1K, where we adopt stronger architectures and optimized training strategies. Finally, we test PLD on MS-COCO for object detection to examine generalization beyond classification.
5.1 CIFAR-100 Classification
We compare PLD with classical and feature-based knowledge distillation methods on CIFAR-100 cifar100 . The dataset contains 50,000 training images and 10,000 validation images across 100 categories at $32\times 32$ resolution. All models are trained from scratch for 250 epochs. We use the AdamW optimizer with $\beta_{1}{=}0.9$ and $\beta_{2}{=}0.999$. The learning rate follows a cosine-annealing schedule, starting at 0.001. We set weight decay to 0.5 and the batch size to 128.
Table 1: Evaluation results on the CIFAR-100 dataset. The upper and lower models denote the teacher and student, respectively.
Method Homogeneous setup Heterogeneous setup
[…]
Feature-based methods
[…]
Logits-based methods
[…]
We apply standard data augmentations: random cropping, horizontal flipping, and per-channel normalization. We report the mean and standard deviation over three independent runs. For DIST huang2022knowledge , we set $\alpha{=}1.0$ for the cross-entropy term, $\beta{=}2.0$ for the inter-class relation loss, $\gamma{=}2.0$ for the intra-class relation loss, and use a temperature of $\tau{=}4.0$. PLD uses the same temperature ($\tau{=}4.0$).
5.2 ImageNet-1K Classification
@@
All experiments on ImageNet-1K deng2009imagenet are conducted following the recipe wightman2021resnet with minimal data augmentation. We train for 100 epochs with an effective batch size of 2048 images (256 per GPU across eight NVIDIA A100 SXM4 80 GB accelerators) using the LAMByou2019large optimizer, an initial learning rate of $5\times 10^{-3}$ decayed by a cosine schedule and linearly warmed up over the first 5 epochs, and weight decay fixed at 0.02. Under this setup, 100 epochs require approximately 6 hours for convolutional backbones and 8 hours for Vision Transformers. To isolate the effect of our losses, we use only random resized crops, horizontal flips, and per-channel normalization.
@@
For validation we adopt the "A-recipe" from wightman2021resnet . Specifically, we set the test resolution $r=224$ and test crop ratio $\rho=0.95$, then $\text{resize}_{\min}=\bigl\lceil r/\rho\bigr\rceil\approx 236,$ apply a bicubic resize of the shorter side to $\text{resize}_{\min}$, followed by a center crop of size $r\times r$. We normalize by the same mean and standard deviation as in training.
@@
Table 2: Top-1 accuracy (%) and student model size across student–teacher pairs on ImageNet-1K. "Teacher Acc" is the teacher’s standalone Top-1; $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ are PLD’s gains over DIST and KD.
@@
Table 3: Top-1 % accuracy of ResNet-50 distilled from diverse teachers on ImageNet-1K. "Params" is the teacher’s parameter count in millions. $\Delta_{\mathrm{DIST}}$ and $\Delta_{\mathrm{KD}}$ report PLD’s gains over DIST and KD, respectively.
@@
Table 4: Top-1 accuracy (%) of the MobileNet-v4 Medium student distilled from a MobileNet-v4 Large teacher after 100 and 300 training epochs. The $\Delta$" row shows the improvement from 100 to 300 epochs.
@@
5.2.1 Baseline Configuration and Ablation
We evaluate four losses under identical training: CE, KD hinton2015distilling , DIST huang2022knowledge , and PLD. We use a ResNet-50 student and a ResNet-152 teacher He2015 . Each objective replaces only the logit-based term and has comparable cost. Thus, KD (divergence-based) and DIST (correlation-based) are natural baselines for our ranking-based PLD. We sweep the key hyperparameters for each method: $\alpha$ for KD, $\beta,\gamma$ for DIST, and $\tau_{T}$ for PLD. Figure 1(a) shows the effect of the CE mixing weight. Reducing $\alpha$ improves KD and DIST up to a region near $\alpha\approx 0.1$. Setting $\alpha=0$ (no CE) degrades performance. Table 5 reports results with and without CE (see Table 8 for the full ablation). Table 6 lists PLD-specific sweeps.
5.2.2 Ablation of PLD
As PLD weights each position by $\alpha_{k}\;=\;\frac{\exp\bigl(t_{\pi^{*}_{k}}/\tau_{T}\bigr)}{\sum_{j=1}^{C}\exp\bigl(t_{\pi^{*}_{j}}/\tau_{T}\bigr)}$, we sweep $\tau_{T}\in\{0.5,1.0,1.5,2.0,4.0\}$ to assess sensitivity: lower $\tau_{T}$ sharpens the distribution, while higher $\tau_{T}$ flattens the distribution. We also test two special cases: Uniform (ListMLE): $\alpha_{k}=1/C$. Position-aware (p-ListMLE): $\alpha_{k}=q^{T}_{\pi^{*}_{k}}/\sum_{j}q^{T}_{\pi^{*}_{j}}$.
As Table 6 shows, the unsoftened PLD ($\tau_{T}=1$) consistently achieves the highest Top-1/Top-5 accuracies, with uniform ListMLE trailing by 2-3 pp. Overall, the best results occur at $\tau_{T}=1.0$ (no softening) or with only slight perturbations ($1.0\pm\epsilon$), demonstrating that PLD is robust to the exact choice of $\alpha$ distribution and does not require aggressive sharpening or over-softening.
Table 5: Top‐1 and Top‐5 accuracy for baseline methods (ResNet-50 student, ResNet-152 teacher with Top-1 Acc. 79.61%). "–" indicates not applicable. CE: standard cross‐entropy. LS:szegedy2016rethinking cross‐entropy with label smoothing ($\epsilon=0.1$). KD: vanilla distillation with mixing weight $\alpha$ and temperature $\tau=2$. DIST: relational distillation with inter-class weight $\beta$, intra-class weight $\gamma$, and temperature $\tau=1$.
Method $\alpha$ $\beta$ $\gamma$ $\tau$ Top-1 Acc.% Top-5 Acc.%
Teacher (ResNet-152) 79.61
CE 71.35
LS ($\epsilon=0.1$) 73.92
KD 0.00 2.00 75.92 92.82
0.10 2.00 76.80 93.16
DIST 0.00 0.50 0.50 1.00 76.47 93.19
0.10 0.45 0.45 1.00 76.60 93.30
Table 6: Top‐1 and Top‐5 accuracy for PLD variants (ResNet‐50 student, ResNet‐152 teacher with Top‐1 Acc. 79.61%). "–" indicates not applicable.
Variant $\tau$ Top-1 Acc.% Top-5 Acc.%
PLD 0.50 76.06 92.39
1.00 77.30 93.28
1.50 77.17 92.74
2.00 76.46 92.34
4.00 75.22 91.34
PLD–ListMLE 74.11 90.60
PLD–pListMLE 76.60 93.23
5.2.3 Distillation Across Homogeneous Architectures
Figure 2(a) evaluates three backbone families in a homogeneous setting: ResNet He2015 , MobileNet-v4 qin2024mobilenetv4 , and Vision Transformers dosovitskiy2020image . Across convolutional models, PLD outperforms KD by 0.5–2.7 pp and DIST by 0.2–0.7 pp. On ViT, PLD yields +0.72 pp over DIST and +0.30 pp over KD. Averaged over all pairs, PLD improves Top-1 accuracy by +0.42 pp versus DIST and +1.04 pp versus KD. Table 2 reports full Top-1 results for KD, DIST, and PLD.
5.2.4 Distillation Across Diverse Teachers
[…]
5.2.5 Consistency under Extended Training
Our homogeneous experiments (Sec. 5.2.3) show that KD sometimes beats DIST and vice versa. We ask whether PLD consistently surpasses both under longer training. We train a MobileNet-v4 Medium student from a MobileNet-v4 Large teacher for 100 and 300 epochs with identical hyperparameters. Table 4 reports Top-1 on ImageNet-1K, and Figure 1(b) provides a detailed comparison. At 100 epochs, PLD achieves 75.72%. This is +0.32 pp over DIST and +0.86 pp over KD. At 300 epochs, PLD reaches 76.94% (+0.38 pp over DIST; +0.80 pp over KD). The PLD gains with longer training are +1.22 pp, which is nearly twice the standard pretraining gains (+0.67 pp). This scaling is comparable across KD (+1.28 pp) and DIST (+1.16 pp). Thus, PLD preserves and slightly widens its advantage under extended training.
Table 7: Object detection performance on MS-COCO. We distill Faster R-CNN detectors with FPN lin2017feature using DKD zhao2022decoupled and PLD. All models are trained under identical schedules. Results are reported on val2017.
Teacher $\rightarrow$ Student Method AP AP50 AP75 APs APm APl
ResNet-50 $\rightarrow$ MobileNetV2-FPN DKD 29.24 50.42 29.87 16.15 31.11 38.07
ResNet-50 $\rightarrow$ MobileNetV2-FPN PLD 28.91 49.63 29.88 16.05 30.41 38.59
ResNet-101 $\rightarrow$ ResNet-18-FPN DKD 32.11 53.43 33.90 18.46 34.43 41.50
ResNet-101 $\rightarrow$ ResNet-18-FPN PLD 32.47 53.83 34.17 18.50 34.97 42.12
ResNet-101 $\rightarrow$ ResNet-50-FPN DKD 36.54 58.57 39.46 21.74 39.80 47.31
ResNet-101 $\rightarrow$ ResNet-50-FPN PLD 36.60 58.28 39.58 21.37 39.76 47.24
5.3 Object Detection on MS-COCO
PLD extends to tasks where the teacher and student share the same output logits, including object detection and semantic segmentation. We distill Faster R-CNN lin2017feature detectors on MS-COCO coco with standard FPNs. We evaluate two convolutional teacher–student pairs: ResNet-50 $\rightarrow$ MobileNetV2-FPN and ResNet-101 $\rightarrow$ {ResNet-18-FPN, ResNet-50-FPN}. All models use identical hyperparameters, and schedules. We compare PLD to the strong DKD baseline zhao2022decoupled .
Table 7 shows that PLD achieves comparable or slightly better performance than DKD across multiple pairs. PLD improves AP and AP75 in most settings while maintaining similar APs for small objects. These results indicate that PLD transfers structured knowledge for dense prediction without modifying the detector.
@@ 6 Conclusion and Limitation
In this work, we introduced Plackett–Luce Distillation (PLD), a unified, choice-theoretic framework for logit-based knowledge distillation. Empirically, across CIFAR-100, ImageNet-1K, and MS-COCO, PLD achieves consistent gains across diverse architectures and distillation objectives, including divergence-based, correlation-based, and feature-based methods, in both homogeneous and heterogeneous teacher–student pairs. These results show that transferring the teacher’s structured preferences yields stronger and more stable student performance than marginal- or correlation-based objectives.
PLD applies broadly to any setting where the student and teacher share the same logit dimensionality. However, it assumes fully aligned class vocabularies and does not yet support mismatched label sets or incremental class addition. PLD also requires sorting over the $C$ output logits to extract the teacher-optimal permutation, which has $O(C\log C)$ complexity per example. Although efficient in distributed implementations, this cost is higher than the $O(C)$ complexity of standard KD or DIST. Empirical runtime analysis (Appendix I) shows that the overhead is negligible for common architectures. As with other distillation methods, PLD’s benefits depend on teacher confidence and may diminish when the teacher’s softmax distribution is nearly uniform, as illustrated in the loss landscape analysis (Appendix J).
PLD’s choice-theoretic foundation opens several directions for future work. First, adaptive or curriculum-driven weighting schemes could adjust the ranking loss based on sample difficulty. Second, extending PLD to other domains, such as sequence modeling, reinforcement learning, or multitask learning, may offer similar gains. We hope this work encourages further exploration of ranking-based objectives for principled and effective model compression.
Acknowledgments and Disclosure of Funding
This work is partially sponsored by National Key R&D Program of China under Grant 2022YFB3104200, NSFC U24A20235 and 62032003.